# Synthetic Data Reveals Generalization Gaps in Correlated Multiple Instance Learning

### v1

@@ 7 Conclusion
We designed an easy-to-implement synthetic dataset designed to mimic key challenges in MIL for medical imaging, especially the need for context from nearby instances. We then demonstrated the limited ability of conventional MIL on such data, by quantifying the performance gap between the optimal Bayes estimator. More recent correlated MIL methods like TransMIL are still notably worse than optimal even with a labeled dataset of 10000 bags. Very recent methods like smAP come closer, but still fall short.
Outlook. We hope our synthetic dataset enables the development of correlated MIL methods that can be trained effectively with limited labeled data.
Appendix A Handcrafted Parameters
@@
Figure 3: Handcrafted parameters for classification (prediction-aggregation approach).
Figure 4: Handcrafted parameters for classification (prediction-aggregation approach).
Figure 5: Handcrafted parameters for classification (embedding-aggregation approach).
Appendix B Varying Separation
Figure 6: Test AUROC as a function of separation $\Delta$, for data drawn from our shifted-mean MIL binary classification task with $R=3,400$ training samples.
@@ Appendix C Bootstrap Analysis
We use bootstrapping to access the statistical significance of the AUROC difference between the Bayes estimator and smAP for the prediction-aggregation approach trained on 10,000 samples We report the mean and 95% confidence interval over 500 subsamples of the test set.
Figure 7: Bootstrap analysis comparing the AUROC difference between the Bayes estimator and smAP for the prediction-aggregation approach train on 10,000 samples.
Appendix D Distribution over Number of Slices
We have a uniform distribution over $S_{i}$ which can be factored out of the numerator and denominator in the posterior and cancels out
$p(y_{i}=1\mid h_{i})=\frac{p(h_{i}\mid y_{i}=1,S_{i})p(y_{i}=1)\cancel{p(S_{i})}}{p(h_{i}\mid y_{i}=0,S_{i})p(y_{i}=0)\cancel{p(S_{i})}+p(h_{i}\mid y_{i}=1,S_{i})p(y_{i}=1)\cancel{p(S_{i})}}.$ (10)

### v2

@@ 7 Conclusion
We designed an easy-to-implement synthetic dataset designed to mimic key challenges in MIL for medical imaging, especially the need for context from nearby instances. We then demonstrated the limited ability of conventional MIL on such data, by quantifying the performance gap between the optimal Bayes estimator. Correlated MIL methods like TransMIL are still notably worse than optimal even with a labeled dataset of 10000 bags. Very recent methods like smAP come closer, but still fall short.
Outlook. Our work reveals a key gap that must be overcome for MIL to succeed on real medical data. Many popular 3D brain scan datasets where MIL could help contain labeled datasets far smaller than the largest $N$ tested here. For example, RSNA (flanders2020construction), as used in castro2024sm, has only 1150 scans for training and evaluation. We hope our synthetic dataset enables the development of correlated MIL methods that can be trained effectively with limited labeled data and make a difference in disease detection and treatment.
Acknowledgments
The authors gratefully acknowledge support from the Alzheimer’s Drug Discovery Foundation and the U.S. National Institutes of Health (grant # 5R01NS134859-02). MCH is also supported in part by the U.S. National Science Foundation (NSF) via grant IIS # 2338962.
Appendix A Experiments Varying the Class Separation $\Delta$
Here, we study how the prediction quality of various MIL methods varies with the class separation parameter $\Delta$, which controls how similar the discriminative feature values are in the data-generating process for the positive class (mean $\mu+\Delta$, variance $\sigma^{2}$) and negative class (mean $\mu$, variance $\sigma^{2}$).
Experiments here use $R=3$, $N=400$, and $\sigma=1$. Given these settings, the chance of a positive discriminative feature value landing closer to the negative mean ($\mu$) than the positive mean ($\mu+\Delta$) is $Pr(h_{ijm}\leq\mu+\frac{1}{2}\Delta)=\Phi(\frac{(\mu+\frac{1}{2}\Delta)-(\mu+\Delta)}{\sigma})=\Phi(-\frac{1}{2}\Delta)$, where $\Phi$ is the standard Normal CDF. When $\Delta=2$ as in the main paper, this is 0.1587; when $\Delta=3$, this is 0.0068; when $\Delta=4$, this is 0.0228; when $\Delta=5$, this is 0.0062. For $\Delta\geq 4$, any individual discriminative feature has over 97% chance to correctly indicate the class label without much need for its surrounding context.
Indeed, in the actual experimental results below, as $\Delta$ increases, we see all methods improve and approach near perfect classification, except for the Mean pooling baseline.
However, for modest $\Delta$ values like 2 or below, there’s a noticeable gap between the ideal Bayes estimator and the actual method performance. We expect the relatively small gap in AUROC between smAP at the largest $N$ and the Bayes estimator shown in Fig. 1 for $\Delta=2$ would notably increase as $\Delta$ gets smaller: note below a gap of roughly 0.1 between smAP and the Bayes estimator at $\Delta=1$.
Figure A.1: Test AUROC as a function of separation $\Delta$. All data drawn from our Shifted Mean MIL data-generating process fro binary classification, with R=3R{=}3 and N=400N{=}400.
Appendix B Handcrafted Parameters
@@
Figure B.1: Handcrafted parameters (prediction-aggregation approach).
Figure B.2: Handcrafted parameters (embedding-aggregation approach).
Figure B.3: Handcrafted parameters (prediction-aggregation approach).
@@ Appendix C Bootstrap Analysis
We use bootstrapping (foody2009classification) to access the statistical significance of the AUROC difference between the Bayes estimator and smAP for the prediction-aggregation approach trained with N=10000N{=}10000. We report the mean and 95% confidence interval over 500 subsamples of the test set.
Figure C.1: Bootstrap analysis comparing the AUROC difference between the Bayes estimator and smAP for the prediction-aggregation approach trained with N=10000N{=}10000.
Appendix D Details Needed for Derivation of Bayes Estimator
One reviewer asked whether the generative process for the number of instances per bag $S_{i}$ needs to be accounted for in the Bayes estimator derivation. We can show that because this distribution is uniform and not dependent on class label, that is $p(S_{i},y_{i})=p(y_{i})p(S_{i})$, the relevant term can be factored out of the numerator and denominator in the label-given-features posterior and canceled
$p(y_{i}=1\mid h_{i},S_{i})=\frac{p(h_{i}\mid y_{i}=1,S_{i})p(y_{i}=1)\cancel{p(S_{i})}}{p(h_{i}\mid y_{i}=0,S_{i})p(y_{i}=0)\cancel{p(S_{i})}+p(h_{i}\mid y_{i}=1,S_{i})p(y_{i}=1)\cancel{p(S_{i})}}.$ (9)
For brevity, we omitted the dependence on $S_{i}$ in many statements in the main paper. The complete class-conditional likelihood PDFs needed to evaluate the right-hand-side above are:
$p(h_{i}|y_{i}=0,S_{i})$ $=\prod_{j=1}^{S_{i}}\prod_{k=1}^{M}\text{NormPDF}(h_{ijk}|\mu,\sigma^{2})$ (10)
$p(h_{i}|y_{i}=1,S_{i})$ $=\sum_{u=1}^{S_{i}-R+1}\left[\underbrace{\frac{1}{S_{i}{-}R{+}1}}_{p(u|y_{i}=1,S_{i})}\prod_{j=1}^{S_{i}}\prod_{k=1}^{M}\underbrace{\text{NormPDF}(h_{ijk}|\mu+\Delta\delta_{u,j,k},\sigma^{2})}_{p(h_{ijk}|u,y_{i}=1,S_{i})}\right]$ (11)
where $\delta_{u,j,k}$ is 1 if $j\in[u,u+R-1]$ and $k$ is a discriminative feature, and 0 otherwise.